OpenStudy (anonymous):

A 0.145 kg baseball is thrown from a platform above the ground at a 10 m/s. The ball is moving 15 m/s when it strikes the ground. From what height above the ground was the ball thrown?

OpenStudy (anonymous):

@Loser66 Here is the new question :)

OpenStudy (loser66):

find t from v = v_0 + at where a = g = 9.8 m/s^2 then plug t into the formula h = h_0 + v0t + 1/2 at^2 where h0 is the height we need to find and h =0 ( when hitting the ground, the ball is get 0 height)

OpenStudy (anonymous):

So the first part is 10 * 9.8 * t = 15

OpenStudy (anonymous):

So t = 0.15?

OpenStudy (anonymous):

Is clear what the initial direction of the ball was? We may need vector addition here.

OpenStudy (anonymous):

I don't think so... But I guess it's possible...

OpenStudy (anonymous):

@wolfe8 , could you help me finish? Sorry I keep asking for your help :P Hope I'm not a bother.

OpenStudy (wolfe8):

That is supposed to be an addition not multiplication 10 + 9.8 * t = 15

OpenStudy (anonymous):

Okay.. So it would really be -4?

OpenStudy (anonymous):

Actually -5 makes more sense :P Sorry

OpenStudy (wolfe8):

I guess we assume it is thrown horizontally. No you cannot have negative time :P See, first you subtract both sides with 10 then divide with 9.8

OpenStudy (anonymous):

OH DUH! Silly me! I've been doing to much math today haha My brain is fried!

OpenStudy (anonymous):

0.51 = t :P

OpenStudy (anonymous):

\[0 = h _{i} + 10 * 0.51 + \frac{ 1 }{ 2 } * 9.8 * 0.51^2\]

OpenStudy (anonymous):

Does that look okay?

OpenStudy (anonymous):

\[h _{i} = -6.4?\]

OpenStudy (wolfe8):

Yup it will be the magnitude so 6.4 m. We get negative because we took acceleration downward (g) as positive in the calculations. Am I right, @Loser66 ?

OpenStudy (anonymous):

Loser66 signed off earlier, but I'm glad you agree! So my answer would be about 6 m?

OpenStudy (wolfe8):

Wait, or maybe we have to use -g in that.

OpenStudy (anonymous):

Hm... That would give an answer of about 4, correct?

OpenStudy (mrnood):

@wolfe8 - If the initial velocity is horizontal then it does not affect the vertical velocity. As Douglas says - there may be a vector question here. In the absence of info I personally would have assumed that the initial velocity was vertically downwards. v = u+at solve for t then s = ut +0.5at^2 u and g and s have the same direction so solve for s. s is the distance below the point thrown.

OpenStudy (wolfe8):

HMMMMMMM. I am getting confused. Well since we used positive for both equations, I think we just go with positive 6.4 m. Yea that's what I was thinking but I decided to just work with what Loser66 said. @MrNood

OpenStudy (anonymous):

Huh, maybe @MrNood is right

OpenStudy (anonymous):

Which is which variable, again? I'm getting confused with all of these numbers and letters O.o g is gravity... S is the distance below the point thrown... what is u?

OpenStudy (mrnood):

v= final velocity u = initial velocity s = displacement (distance) t=time

OpenStudy (anonymous):

So t is still 0.51

OpenStudy (mrnood):

yes

OpenStudy (wolfe8):

Yea they're the same equations and what Loser said is the same thing. I just didn't realize it.

OpenStudy (anonymous):

So my answer is about 6.3

OpenStudy (anonymous):

Okay, so it is still the same! Good! I was scared for a minute there

OpenStudy (anonymous):

Thank you for your help everyone!!!

OpenStudy (wolfe8):

You're welcome. I just helped with the math btw. I went with Loser's physics. And MrNood confirmed it.

OpenStudy (anonymous):

You were all a big help :P! I needed all three of you! haha :)